# Imprints of the Local Bubble and Dust Complexity on Polarized Dust Emission

### v1

@@ 5 Discussion and Conclusions
Given the significance of the 3D dust distribution on measurements of the polarized dust emission, these data can be combined with position-position-velocity maps of the neutral hydrogen-based dust polarization templates. These templates, constructed based on the orientation of neutral hydrogen filaments, have been shown to correlate very well with the measured dust polarization (Clark & Hensley, 2019; Cukierman et al., 2023; BICEP/Keck Collaboration et al., 2023; Halal et al., 2024). Since the neutral hydrogen and dust trace similar volumes of the diffuse interstellar medium (Boulanger et al., 1996; Lenz et al., 2017), future work could morphologically match the position-position-position dust maps with the position-position-velocity neutral hydrogen-based maps to form 4D position-position-position-velocity maps of the magnetic field and polarized dust emission. Starlight polarization can also be used to provide a tomographic view of the plane-of-the-sky magnetic field and polarized dust emission for sightlines with these measurements (Panopoulou et al., 2019; Tassis et al., 2018). These data combined with Faraday tomography (Van Eck et al., 2017) or rotation measures (Tahani et al., 2018) can be used to constrain the 3D magnetic field structure.
@@
We thank Gordian Edenhofer and Minjie Lei for insightful discussions. This work was supported by the National Science Foundation under grant No. 2106607. MT is supported by the Banting Fellowship (Natural Sciences and Engineering Research Council Canada) hosted at Stanford University and the Kavli Institute for Particle Astrophysics and Cosmology (KIPAC) Fellowship. This publication utilizes data from Planck, an ESA science mission funded by ESA Member States, NASA, and Canada. The computations in this paper were run on the Sherlock cluster, supported by the Stanford Research Computing Center at Stanford University.

### v2

@@ 5 Discussion and Conclusions
Given the significance of the 3D dust distribution on measurements of the polarized dust emission, these data can be combined with position-position-velocity maps of the neutral hydrogen-based dust polarization templates. These templates, constructed based on the orientation of neutral hydrogen filaments, have been shown to correlate very well with the measured dust polarization (Clark & Hensley, 2019; Cukierman et al., 2023; BICEP/Keck Collaboration et al., 2023; Halal et al., 2024). Since the neutral hydrogen and dust trace similar volumes of the diffuse interstellar medium (Boulanger et al., 1996; Lenz et al., 2017), future work could morphologically match the position-position-position dust maps with the position-position-velocity neutral hydrogen-based maps to form 4D position-position-position-velocity maps of the magnetic field and polarized dust emission. Starlight polarization can also be used to provide a tomographic view of the plane-of-the-sky magnetic field and polarized dust emission for sightlines with these measurements (Panopoulou et al., 2019; Tassis et al., 2018; Pelgrims et al., 2023, 2024). These data combined with Faraday tomography (Van Eck et al., 2017) or rotation measures (Tahani et al., 2018) can be used to constrain the 3D magnetic field structure.
@@
We thank Gordian Edenhofer and Minjie Lei for insightful discussions. This work was supported by the National Science Foundation under grant No. AST-2106607. MT is supported by the Banting Fellowship (Natural Sciences and Engineering Research Council Canada) hosted at Stanford University and the Kavli Institute for Particle Astrophysics and Cosmology (KIPAC) Fellowship. This publication utilizes data from Planck, an ESA science mission funded by ESA Member States, NASA, and Canada. The computations in this paper were run on the Sherlock cluster, supported by the Stanford Research Computing Center at Stanford University.
Appendix A Analysis Variations
$\Delta_{n}$$\#_{\rm voxels}$ low complexity high complexity ##\## of matched pairs $\mu_{\Delta p_{353}}$ $\sigma_{\Delta p_{353}}$ p-value
$\mathbf{7\times 10^{-6}}$ 3 $\mathbf{\mathcal{N}_{c}\leq 1.1}$ $\mathbf{\mathcal{N}_{c}\geq 1.5}$ 753 $\mathbf{-1.47\times 10^{-2}}$ $\mathbf{0.22\times 10^{-2}}$ $\mathbf{<0.001}$
$7\times 10^{-6}$ 3 $\mathcal{N}_{c}\leq 1.4$ $\mathcal{N}_{c}\geq 2.3$ 4,674 $-1.10\times 10^{-2}$ $0.08\times 10^{-2}$ $<0.001$
$9\times 10^{-6}$ 3 $\mathcal{N}_{c}\leq 1.1$ $\mathcal{N}_{c}\geq 1.5$ 776 $-1.33\times 10^{-2}$ $0.26\times 10^{-2}$ $<0.001$
$9\times 10^{-6}$ 3 $\mathcal{N}_{c}\leq 1.4$ $\mathcal{N}_{c}\geq 2.3$ 4,719 $-1.01\times 10^{-2}$ $0.12\times 10^{-2}$ $<0.001$
$5\times 10^{-6}$ 5 $\mathcal{N}_{c}\leq 1.1$ $\mathcal{N}_{c}\geq 1.5$ 1,005 $-0.98\times 10^{-2}$ $0.20\times 10^{-2}$ $<0.001$
$5\times 10^{-6}$ 5 $\mathcal{N}_{c}\leq 1.4$ $\mathcal{N}_{c}\geq 2.3$ 5,311 $-0.82\times 10^{-2}$ $0.14\times 10^{-2}$ $<0.001$
$3\times 10^{-6}$ 5 $\mathcal{N}_{c}\leq 1.1$ $\mathcal{N}_{c}\geq 1.5$ 984 $-1.00\times 10^{-2}$ $0.21\times 10^{-2}$ $<0.001$
$3\times 10^{-6}$ 5 $\mathcal{N}_{c}\leq 1.4$ $\mathcal{N}_{c}\geq 2.3$ 5,284 $-0.80\times 10^{-2}$ $0.13\times 10^{-2}$ $<0.001$
Table 1: Results obtained from varying both the dendrogram parameters ($\Delta_{n}$ and #$\#_{\rm voxels}$) used to identify the density peaks in the line-of-sight dust distribution and from varying the upper and lower thresholds in $\mathcal{N}_{c}$ used to divide the sightlines into low- and high-complexity groups. The results of the main analysis are bolded in the first row. The mean and standard deviation in the $\mu_{\Delta p_{353}}$ and $\sigma_{\Delta p_{353}}$ columns are over the 12 posterior samples of the Edenhofer et al. (2023) maps.
$\Delta_{n}$$\#_{\rm voxels}$ low complexity high complexity ##\## of matched pairs $\mu_{\Delta p_{353}}$ $\sigma_{\Delta p_{353}}$ p-value
$7\times 10^{-6}$ 3 $\mathcal{N}_{c}\leq 1.1$ $\mathcal{N}_{c}\geq 1.5$ 753 $-1.45\times 10^{-2}$ $0.23\times 10^{-2}$ $<0.001$
$7\times 10^{-6}$ 3 $\mathcal{N}_{c}\leq 1.4$ $\mathcal{N}_{c}\geq 2.3$ 4,674 $-1.02\times 10^{-2}$ $0.08\times 10^{-2}$ $<0.001$
$9\times 10^{-6}$ 3 $\mathcal{N}_{c}\leq 1.1$ $\mathcal{N}_{c}\geq 1.5$ 776 $-1.31\times 10^{-2}$ $0.34\times 10^{-2}$ $<0.001$
$9\times 10^{-6}$ 3 $\mathcal{N}_{c}\leq 1.4$ $\mathcal{N}_{c}\geq 2.3$ 4,719 $-1.00\times 10^{-2}$ $0.15\times 10^{-2}$ $<0.001$
$5\times 10^{-6}$ 5 $\mathcal{N}_{c}\leq 1.1$ $\mathcal{N}_{c}\geq 1.5$ 1,005 $-0.60\times 10^{-2}$ $0.17\times 10^{-2}$ $<0.001$
$5\times 10^{-6}$ 5 $\mathcal{N}_{c}\leq 1.4$ $\mathcal{N}_{c}\geq 2.3$ 5,311 $-0.82\times 10^{-2}$ $0.13\times 10^{-2}$ $<0.001$
$3\times 10^{-6}$ 5 $\mathcal{N}_{c}\leq 1.1$ $\mathcal{N}_{c}\geq 1.5$ 984 $-0.67\times 10^{-2}$ $0.22\times 10^{-2}$ $<0.001$
$3\times 10^{-6}$ 5 $\mathcal{N}_{c}\leq 1.4$ $\mathcal{N}_{c}\geq 2.3$ 5,284 $-0.80\times 10^{-2}$ $0.13\times 10^{-2}$ $<0.001$
Table 2: Results obtained from the same variations described in Table 1, where the sightline matching over the two complexity groups in this case includes the absolute value of the Galactic latitude as an additional confounding variable. The mean and standard deviation in the $\mu_{\Delta p_{353}}$ and $\sigma_{\Delta p_{353}}$ columns are over the 12 posterior samples of the Edenhofer et al. (2023) maps.
$\Delta_{n}$$\#_{\rm voxels}$ $\mu_{\Delta p_{353}}$ $\sigma_{\Delta p_{353}}$ p-value
$7\times 10^{-6}$ 3 $-1.08\times 10^{-2}$ $0.17\times 10^{-2}$ $<0.001$
$9\times 10^{-6}$ 3 $-1.14\times 10^{-2}$ $0.24\times 10^{-2}$ $<0.001$
$5\times 10^{-6}$ 5 $-0.84\times 10^{-2}$ $0.17\times 10^{-2}$ $<0.001$
$3\times 10^{-6}$ 5 $-0.89\times 10^{-2}$ $0.18\times 10^{-2}$ $<0.001$
Table 3: Results obtained from sampling 1,000 sightlines from each of the $\mathcal{N}_{c}~{}\leq~{}1.4$ and $\mathcal{N}_{c}~{}\geq~{}2.3$ complexity groups before matching. The different rows are for different variations in the dendrogram parameters ($\Delta_{n}$ and #$\#_{\rm voxels}$) used in identifying the density peaks in the line-of-sight dust distribution. The mean and standard deviation in the $\mu_{\Delta p_{353}}$ and $\sigma_{\Delta p_{353}}$ columns are over the 12 posterior samples of the Edenhofer et al. (2023) maps.
In this appendix, we list some of the results obtained from varying the main analysis choices in Section 4. These analysis variations are described and their results summarized in Section 4.6. In Table 1, we present some of the results obtained from varying the dendrogram parameters used in identifying the density peaks in the line-of-sight dust distribution and from varying the upper and lower thresholds in $\mathcal{N}_{c}$ used for splitting the sightlines into low- and high-complexity groups. Table 2 lists the results of the same variations performed in Table 1 but when including the absolute value of the Galactic latitude as an additional confounding variable in the sightline matching across the two complexity groups. Finally, we present the results of sampling 1,000 sightlines from each of the two complexity groups with $\mathcal{N}_{c}\leq 1.4$ and $\mathcal{N}_{c}\leq 2.3$ before matching in Table 3. We find that the result in our main analysis is robust to all of these analysis variations.